4a+3c=-2a+5
Solve for A
And the Answer has to like this
a=(___)/____

Answer by jim_thompson5910(35256)

4a+3c=-2a+5

4a+3c+2a=5

4a+2a=5-3c

6a = 5-3c

a = (5-3c)/6